OpenStudy (anonymous):

Check my answer please! Eliminate the parameter x = t + 4, y = t^2

OpenStudy (anonymous):

I got y = squareroot x - 4

OpenStudy (anonymous):

@satellite73 Did I do this right?

OpenStudy (agent0smith):

How did squared become square root??

OpenStudy (agent0smith):

Show your work because you're almost right

OpenStudy (anonymous):

Well don't I have to square root the side to get t alone?

OpenStudy (agent0smith):

Show your work

OpenStudy (anonymous):

x= t + 4 -4 x-4 = t y=t^2 squareroot y = t

OpenStudy (anonymous):

hmm.. kinda looks funny now.

OpenStudy (agent0smith):

You've done nothing wrong (in that working). Your final answer earlier, was wrong.

OpenStudy (anonymous):

y = squareroot x + x + 4?

OpenStudy (agent0smith):

You're making this more complicated than it is. You have: x-4 = t and: y = t^2 You want: y = ...

OpenStudy (anonymous):

y= (x-4)^2 y= x^2 + 16?

OpenStudy (anonymous):

oops

OpenStudy (agent0smith):

y= (x-4)^2 <=== yes y= x^2 + 16 <=== no, do not make this mistake... you can't just distribute the exponent like that.

OpenStudy (anonymous):

x6@ -8x + 16

OpenStudy (anonymous):

y=x^2 - 8x + 16